# YOLO-LLTS: Real-Time Low-Light Traffic Sign Detection via Prior-Guided Enhancement and Multi-Branch Feature Interaction

### v3

@@ IV-D3 Effectiveness of MFIA
IV-E Error Analysis and Model Deployment
In this study, we conducted a detailed analysis of the training errors of the traffic sign detection model across three different datasets [70]. Fig. 14 presents the evolution of key training metrics, including normalized loss, precision, recall, mAP50, and mAP50-95, with respect to the number of training epochs. The normalized loss is the sum of the box loss, classification loss, and distribution focal Loss, which we have normalized for ease of presentation. As shown in the figure, the normalized loss decreases rapidly and eventually stabilizes as the number of training epochs increases across all datasets, indicating that the model exhibits good convergence on these datasets. These critical performance metrics quickly increase in the early stages of training and eventually stabilize.
At the same time, we test the detection accuracy and inference times on mobile edge devices, NVIDIA Jetson AGX Xavier and NVIDIA Jetson AGX Orin. As shown in Fig. 15 and Fig. 16, we first connect the device to the display, keyboard, and mouse, then connect the network cable and power supply. We power on the edge computing device and test the trained weights on both the Jetson AGX Xavier and Jetson AGX Orin. The experiments on Jetson AGX Xavier were conducted in an Ubuntu 20.04 operating system environment, with optimization support provided by Jetpack 5.1. The experiments on Jetson AGX Orin were conducted in an Ubuntu 18.04 operating system environment, using the PyTorch deep learning framework (version 2.1.0), with optimization support provided by Jetpack 5.1. We selected 500 nighttime images from the CCTSDB2021 dataset for testing. As shown in Tab. VII and Tab. VIII, YOLO-LLTS achieves the best performance in terms of accuracy. Specifically, our model outperforms the average of YOLOv8 to YOLOv12 by 4.6 in accuracy, 6.7 in recall, and 5.8 in F1 score. It also exceeds the average by 7.4 in mAP50 and 4.7 in mAP50:95. Without utilizing TensorRT acceleration, the number of parameters on the Xavier is 10.2 million, with an inference time of 138.8 ms. On the Orin, the number of parameters is 10.2 million, with an inference time of 44.9 ms. The experiment demonstrates that our model maintains relatively competitive performance even on devices with constrained power.
@@ V Conclusion
Under low-light conditions, the perception systems of autonomous vehicles are prone to false positives and missed detections, leading to road traffic accidents. In this paper, we propose YOLO-LLTS, an end-to-end real-time traffic sign detection algorithm, which effectively enhances the accuracy and speed of traffic sign recognition, thereby improving the safety of ADAS and autonomous vehicles in low-light environments. To address the lack of nighttime-specific traffic sign datasets, we constructed a novel dataset named Chinese Nighttime Traffic Sign Sample Set (CNTSSS). This dataset includes images captured under varying low-light conditions from dusk to midnight, covering diverse scenarios such as urban, highway, and rural environments under different weather conditions. We introduce the High-Resolution Feature Map for Small Object Detection (HRFM-SOD) module to effectively address the indistinct features of small objects under low-light conditions, significantly improving both detection accuracy and inference speed. Moreover, we design the Multi-branch Feature Interaction Attention (MFIA) module, enabling deep interaction and fusion of features across multiple receptive fields, thereby enhancing the model’s ability to capture and utilize critical information. Furthermore, we propose the Prior-Guided Feature Enhancement Module (PGFE) to alleviate challenges such as increased noise, reduced contrast, and blurriness under low-light environments, significantly boosting the detection performance. Experimental results demonstrate that our approach achieves state-of-the-art performance on the TT100K-night, CNTSSS, CCTSDB2021, and GTSDB-night datasets. The deployment experiments on edge devices further validate the practical effectiveness and real-time applicability of our method.
In future work, we plan to expand our CNTSSS dataset with more diverse scenarios. In addition, although the YOLO-LLTS algorithm shows significant improvements in traffic sign detection and recognition under low-light conditions, it still has limitations in extremely low-light conditions. As shown in Fig. 18, when the road environment has weak lighting or lacks street lamps, our model failed to recognize traffic signs. Future research could further explore ways to enhance the model’s detection performance in extremely low-light or nighttime environments.

### v4

@@ V Conclusion
Under low-light conditions, the perception systems of autonomous vehicles are prone to false positives and missed detections, leading to road traffic accidents. In this paper, we propose YOLO-LLTS, an end-to-end real-time traffic sign detection algorithm, which effectively enhances the accuracy and speed of traffic sign recognition, thereby improving the safety of ADAS and autonomous vehicles in low-light environments. To address the lack of nighttime-specific traffic sign datasets, we constructed a novel dataset named Chinese Nighttime Traffic Sign Sample Set (CNTSSS). This dataset includes images captured under varying low-light conditions from dusk to midnight, covering diverse scenarios such as urban, highway, and rural environments under different weather conditions. We introduce the High-Resolution Feature Map for Small Object Detection (HRFM-SOD) module to effectively address the indistinct features of small objects under low-light conditions, significantly improving both detection accuracy and inference speed. Moreover, we design the Multibranch Feature Interaction Attention (MFIA) module, enabling deep interaction and fusion of features across multiple receptive fields, thereby enhancing the model’s ability to capture and utilize critical information. Furthermore, we propose the Prior-Guided Feature Enhancement Module (PGFE) to alleviate challenges such as increased noise, reduced contrast, and blurriness under low-light environments, significantly boosting the detection performance. Experimental results demonstrate that our approach achieves state-of-the-art performance on the TT100K-night, CNTSSS, CCTSDB2021, and GTSDB-night datasets. The deployment experiments on edge devices further validate the practical effectiveness and real-time applicability of our method.
Although the YOLO-LLTS algorithm shows significant improvements in traffic sign detection and recognition under low-light conditions, there are still some failure cases. As shown in Fig. 20(a), when the road environment has weak lighting or lacks street lamps, our model encounters false positives and false negatives. As shown in Fig. 20(b), when the target is extremely small, the model suffers from missed detections. As shown in Fig. 20(c), when the sign is obstructed by barriers, the model fails to detect the target. Future research could further explore ways to enhance the model’s detection performance under extremely low light, extremely small objects, and partially occluded objects conditions. Additionally, both our CNTSSS dataset and existing public datasets do not contain data on extreme weather conditions, such as heavy fog or heavy rain. Therefore, the model’s effectiveness under extreme weather remains unverified, posing a risk of false negatives and false positives. In the future, we plan to expand the CNTSSS dataset to include data on extreme weather conditions and optimize the model to better adapt to challenging weather situations. Furthermore, we plan to expand the CNTSSS dataset to include more countries or regions, systematically assessing the model’s robustness across different traffic regulatory environments. The code and the dataset are available at https://github.com/linzy88/YOLO-LLTS.